OpenStudy (anonymous):

6x^5+12x^1 ------------- -2x^-3

OpenStudy (campbell_st):

the problem can be simplifed by splitting into 2 fractions \[\frac{6x^5}{-2x^{-3} }+ \frac{12x^1}{-2x^{-3}}\] so you need to simplify 6/-2 and 12/-2 then the index law for division is subtract the powers \[\frac{x^a}{x^b} = x^{a - b}\] if you subtract the powers, you'll get the correct power of x. But be careful as you'll be subtracting a negative...